OpenStudy (goten77):

if i said differentiate y=2x does that mean take the derivative? or what exactly?

OpenStudy (anonymous):

yes

OpenStudy (goten77):

well what is it then when you get dy by it self? and have something like dy = 2dx

OpenStudy (abb0t):

separation of variables. take the integral of each.

OpenStudy (goten77):

something like the differential?

OpenStudy (unklerhaukus):

the derivative of a straight line is simply the slope

OpenStudy (anonymous):

Differentiate \[y=2x\]\[\frac{ dy }{ dx }=2\]\[dy=2dx\]that looks familiar Integrate:\[\int\limits dy=\int\limits 2 \space dx\]\[y=2x+C\]

OpenStudy (abb0t):

\[\int\limits dy = \int\limits 2xdx \]

OpenStudy (goten77):

so differentiate just means put the dy by it self?

OpenStudy (goten77):

<.< using words that are so similar is not a good idea

OpenStudy (abb0t):

you are basically taking the antiderivative, which in essence is differentiating.